OpenStudy (anonymous):

write an equation for the line parallel to the given line C. C(3,6) y=-2x+3

OpenStudy (anonymous):

the line paralle to the above line is of the form y=-2x +k (just for parallel lines constant term is different as their slopes are same ) since our line passes throug(3,6) we have 6=-2(3) +k or k=12 hence the reqd line is y=-2x +12

OpenStudy (anonymous):

what would that look like in slope intercept form?

OpenStudy (anonymous):

And how did you find k?

jimthompson5910 (jim_thompson5910):

y = mx+b y = -2x+b 6 = -2(3) + b 6 = -6 + b 6+6 = b 12 = b b = 12 So the parallel equation is y = -2x + 12

OpenStudy (anonymous):

thanks!